# Multi-modal Attribute Prompting for Vision-Language Models

### v2

@@ 4 Experiments
Implementation Details. In all the experiments, we use the pre-trained CLIP  [38] with ViT-B/16 image encoder backbone as the base model. We use GPT-3.5 as the LLM. For MAP, we set the number of textual attribute prompts $N$ to 4, and the number of visual attribute prompts $M$ to 4. The AVAE module is inserted into the 7th transformer layer in the Vision Transformer (ViT). The default value of $\lambda$ is set as 10. $\beta$ is set as 1. We train the model using the SGD optimizer with a learning rate of 0.002. For the base-to-novel generalization setting, the model is trained for 20 epochs with a batch size of 16. For few-shot image classification, the maximum epoch is set to 200 for 16/8 shots, 100 for 4/2 shots, and 50 for 1 shot (except for ImageNet, where the maximum epoch is fixed to 50).
4.1 Base-to-Novel Generalization
Table 1: Accuracy comparison on Base-to-novel generalization of MAP with previous methods. HM: Harmonic mean to highlight the generalization trade-off [45].
     Dataset      CLIP      CoOp      CoCoOp      ProDA      VDT-Adapter      MaPLe      MAP
     [38]      [5]      [50]      [29]      [31]      [21]      (Ours)
      Average on 11 datasets      Base      69.34      82.69      80.47      81.56      82.48      82.28      83.66
     Novel      74.22      63.22      71.69      72.30      74.51      75.14      75.76
     HM      71.70      71.66      75.83      76.65      78.09      78.55      79.36
     ImageNet      Base      72.43      76.47      75.98      75.40      76.4      76.66      76.60
     Novel      68.14      67.88      70.43      70.23      68.3      70.54      70.60
     HM      70.22      71.92      73.10      72.72      72.12      73.47      73.48
     Caltech101      Base      96.84      98.00      97.96      98.27      98.3      97.74      98.30
     Novel      94.00      89.81      93.81      93.23      95.9      94.36      93.80
     HM      95.40      93.73      95.84      95.68      97.09      96.02      96.00
     OxfordPets      Base      91.17      93.67      95.20      95.43      94.4      95.43      95.43
     Novel      97.26      95.29      97.69      97.83      97.0      97.76      96.90
     HM      94.12      94.47      96.43      96.62      95.68      96.58      96.16
      Stanford Cars      Base      63.37      78.12      70.49      74.70      76.8      72.94      76.70
     Novel      74.89      60.40      73.59      71.20      72.9      74.00      73.73
     HM      68.65      68.13      72.01      72.91      74.80      73.47      75.18
     Flowers102      Base      72.08      97.60      94.87      97.70      97.4      95.92      97.57
     Novel      77.80      59.67      71.75      68.68      75.3      72.46      75.23
     HM      74.83      74.06      81.71      80.66      84.94      82.56      84.95
     Food101      Base      90.10      88.33      90.70      90.30      90.4      90.71      90.30
     Novel      91.22      82.26      91.29      88.57      91.2      92.05      89.30
     HM      90.66      85.19      90.99      89.43      90.80      91.38      89.80
      FGVC Aircraft      Base      27.19      40.44      33.41      36.90      37.8      37.44      41.63
     Novel      36.29      22.30      23.71      34.13      33.0      35.61      36.43
     HM      31.09      28.75      27.74      35.46      35.24      36.50      38.84
     SUN397      Base      69.36      80.60      79.74      78.67      81.4      80.82      82.33
     Novel      75.35      65.89      76.86      76.93      76.8      78.70      76.30
     HM      72.23      72.51      78.27      77.79      79.03      79.75      79.20
     DTD      Base      53.24      79.44      77.01      80.67      81.8      80.36      82.63
     Novel      59.90      41.18      56.00      56.48      62.3      59.18      66.23
     HM      56.37      54.24      64.85      66.44      70.73      68.16      73.53
     EuroSAT      Base      56.48      92.19      87.49      83.90      88.5      94.07      92.13
     Novel      64.05      54.74      60.04      66.00      70.5      73.23      76.10
     HM      60.03      68.69      71.21      73.88      78.48      82.35      83.33
     UCF101      Base      70.53      84.69      82.33      85.23      84.1      83.00      86.67
     Novel      77.50      56.05      73.45      71.97      76.4      78.66      78.77
     HM      73.85      67.46      77.64      78.04      80.07      80.77      82.52
@@ 4.1 Base-to-Novel Generalization
Table 2: Domain generalization. Prompting methods are trained on ImageNet and evaluated on datasets with domain shifts.
@@
ImageNet ImageNet-V2 ImageNet-S ImageNet-A ImageNet-R Avg.
CoOp 71.51 64.20 47.99 49.71 75.21 59.28
CoCoOp 71.02 64.07 48.75 50.63 76.18 59.91
MaPLe 70.72 64.07 49.15 50.90 76.98 60.27
@@
Table 3: Cross-dataset evaluation. Prompting methods are trained on ImageNet and evaluated on target datasets. MAP achieves overall favorable performance.
@@
CoOp 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55
CoCoOp 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21
MaPLe 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69
@@
Compared to CoOp, MAP exhibits higher harmonic mean accuracy across all datasets. As shown in Table 1, MAP, on average, increases novel accuracy by 12.54% and base accuracy by 0.97%. This demonstrates that MAP not only enhances the model’s generalization to novel classes but also achieves better alignment between visual and textual modalities within base classes.
We also provide results of several recent methods in Table 1, MAP outperforms other methods in both base classes and novel classes. It’s worth noting that VDT-Adapter [31], which utilizes textual attributes obtained from GPT-4 to construct prompts, improves novel accuracy compared to CoOp. However, it neglects modeling visual attributes and fails to fully leverage the role of attributes. MAP outperforms VDT-Adapter 1.18% in base classes and 1.25% in novel classes.
Figure 5: Average few-shot image classification results over 6 datasets. (a) Absolute accuracy improvements provided by using AVAE compared to scenarios without AVAE. (b) The impact of inserting AVAE into different layers of ViT with 1 shot per class.
4.2 Few-Shot Image Classification
To evaluate few-shot learning ability, we adopt the few-shot evaluation protocol from CLIP [38], utilizing 1, 2, 4, 8, and 16 shots per class for training and deploying models in full test sets. Figure 4 summarizes the performance of MAP in few-shot learning on 11 datasets. Each plot compares MAP with CoOp and CoOp+VPT. CoOp+VPT refers to the combination of CoOp and VPT, i.e., the integration of both learnable text prompts and learnable visual prompts [18] into the CLIP model simultaneously. In terms of the overall performance (Figure 4, top-left), compared to CoOp, the combination of CoOp and VPT shows some improvement, though not significant. However, in the 1-shot setting, the performance of the combination is even worse than CoOp alone. This suggests that simply introducing more learnable parameters in the vision encoder brings limited performance improvement in the extreme few-shot setting. However, MAP, consistently delivers significant performance improvements, even in scenarios with very few training samples (e.g., 1-shot), showcasing the effectiveness of our visual attribute prompts enhanced by textual guidance. Furthermore, on certain datasets (Caltech101, Flowers102, DTD, SUN397, and OxfordPets), CoOp+VPT does not outperform CoOp alone, whereas MAP consistently achieves superior performance across all benchmark datasets, demonstrating the generalizability of MAP across diverse datasets.
4.3 Domain Generalization
We evaluate the generalizability of MAP on out-of-distribution data through a comparison with CoOp, CoCoOp, and MaPLe. The overall results are summarized in Table 2. MAP not only attains the highest accuracy on ImageNet but also exhibits superior performance on ImageNetV2, ImageNet-A, and ImageNet-R, demonstrating the robustness of MAP.
4.4 Cross-Dataset Evaluation
Table 3 summarizes the results of MAP and previous methods on cross-dataset evaluation benchmark. On the source dataset, MAP achieves the highest score. Compared with CoOp, CoCoOp, and MaPLe, MAP shows favorable performance and achieves better generalization in 7/10, 6/10, and 6/10 datasets respectively.
4.5 Ablation Study
@@
Effectiveness of Attribute Prompts. We denote Textual Attribute Prompts as TAP and Visual Attribute Prompts as VAP. We remove TAP and VAP from MAP as our baseline. The results in Table 4 are analyzed as follows: (1) Compared to the baseline, utilizing TAP powered by the LLM effectively improves the novel accuracy, achieving an accuracy gain of 1.43%, which demonstrates textual attributes enrich the semantics for novel classes. (2) The incorporation of VAP shows a distinct performance boost on both base (+1.6%) and novel classes (+2.11%). This proves that VAP contributes to enhancing fine-grained visual perception ability by capturing visual attributes.
Table 4: Ablation results averaged over 11 datasets in the base-to-novel setting.
@@
Effectiveness of Adaptive Visual Attribute Enhancement. To verify the accuracy improvement when using AVAE, we conduct few-shot image classification experiments on 6 datasets (Flo, DTD, UCF, Pet, Cal, Foo). As shown in Figure 5 (a), the employment of AVAE brings remarkable performance gains. Furthermore, we investigate the impact of placing AVAE into different ViT layers. As observed from Figure 5 (b), placing AVAE in the middle layers (Layer 6-8) attains superior performance. When applying AVAE in the shallow or deep layers, the performance deteriorates obviously compared to the middle layers. Therefore, the AVAE module should be placed in the middle layers. Initial visual attribute prompts can aggregate visual regional features in shallow layers, and continue to capture visual attributes in the remaining layers after enhancement by AVAE.
@@
Figure 7: The visualization of visual attribute prompts for the category “Moon Orchid” and “Japanese Anemone”. Guided by textual attribute semantics, visual attribute prompts focus on distinctive visual details, such as different leaf shapes.
5 Conclusion

### v3

@@ IV Experiments
Implementation Details. In all the experiments, we use the pre-trained CLIP  [1] with ViT-B/16 image encoder backbone as the base model. We use the GPT-3.5 as the large language model. For MAP, we set the number of textual attribute prompts $N$ to 4, and the number of visual attribute prompts $M$ to 4. The AVAE module is inserted into the 7th transformer layer in the Vision Transformer (ViT). The default value of $\lambda$ is set as 10. $\beta$ is set as 1. We train the model using the SGD optimizer with a learning rate of 0.002. For the base-to-novel generalization setting, the model is trained for 20 epochs with a batch size of 16. For few-shot image classification, the maximum epoch is set to 200 for 16/8 shots, 100 for 4/2 shots, and 50 for 1 shot (except for ImageNet, where the maximum epoch is fixed to 50).
IV-A Base-to-Novel Generalization
TABLE III: Comparing MAP against more methods on the average accuracy over 11 datasets.
Method Base Novel HM
CLIP [1] 69.34 74.22 71.70
CoOp [16] 82.69 63.22 71.66
CoCoOp [18] 80.47 71.69 75.83
ProDA [21] 81.56 72.30 76.65
RPO [22] 81.13 75.00 77.78
VDT-Adapter [26] 82.48 74.51 78.09
MaPLe [23] 82.28 75.14 78.55
MAP 83.66 75.76 79.36
Compared to CoOp, MAP exhibits higher harmonic mean accuracy across all datasets. As shown in Table II, MAP, on average, increases novel accuracy by 12.54% and base accuracy by 0.97%. This demonstrates that MAP not only enhances the model’s generalization to novel classes but also achieves better alignment between visual and textual modalities within base classes.
We present the average accuracy results across 11 datasets for MAP compared with several other methods in Table III. MAP outperforms other methods by a significant margin, demonstrating our superior performance over other methods. It’s worth noting that VDT-Adapter [26], which leverages textual attributes obtained from GPT-4 to formulate prompts, improves the novel accuracy compared to CoOp. However, it neglects modeling visual attributes and fails to leverage the role of attributes fully. MAP outperforms VDT-Adapter 1.18% in base classes and 1.25% in novel classes.
Figure 5: The average few-shot image classification results of more methods across 11 datasets.
IV-B Few-Shot Image Classification
To evaluate few-shot learning ability, we adopt the few-shot evaluation protocol from CLIP [1], utilizing 1, 2, 4, 8, and 16 shots per class for training and deploying models in full test sets. Figure 4 summarizes the performance of MAP in few-shot learning on 11 datasets. Each plot compares MAP with CoOp and CoOp+VPT. CoOp+VPT refers to the combination of CoOp and VPT, i.e., the integration of both learnable text prompts and learnable visual prompts [49] into the CLIP model simultaneously. In terms of the overall performance (Figure 4, top-left), compared to CoOp, the combination of CoOp and VPT shows some improvement, though not significant. However, in the 1-shot setting, the performance of the combination is even worse than CoOp alone. This suggests that simply introducing more learnable parameters in the vision encoder brings limited performance improvement in the extreme few-shot setting. However, MAP consistently delivers significant performance improvements, even in scenarios with very few training samples (e.g., 1-shot), showcasing the effectiveness of our visual attribute prompts enhanced by textual guidance. Furthermore, on certain datasets (Caltech101, Flowers102, DTD, SUN397, and OxfordPets), CoOp+VPT does not outperform CoOp alone, whereas MAP consistently achieves superior performance across all benchmark datasets, demonstrating the generalizability of MAP across diverse datasets.
In Figure 5, we present the performance results of additional methods for few-shot image classification. Tip-adapter-F [78], the fine-tuned version of Tip-adapter, requires fine-tuning on the few-shot training data to update the adapter. The results show that Tip-adapter-F consistently achieves better performance than Tip-adapter and Linear probe CLIP. MaPLe [23] achieves performance comparable to Tip-adapter-F overall. Notably, MAP consistently outperforms both MaPLe [23] and Tip-adapter-F [78] in few-shot image classification across various shot settings, highlighting the effectiveness of our proposed approach.
IV-C Domain Generalization
To evaluate the model’s robustness under domain shifts, we initially train the model using the source dataset, ImageNet [66]. Subsequently, we evaluate its performance on target out-of-distribution datasets, namely ImageNetV2 [76], ImageNet-Sketch [77], ImageNet-A [75] and ImageNet-R [74]. The overall results are summarized in Table IV. From the experimental results, the fully fine-tuned CLIP model shows poorer performance compared to the zero-shot CLIP on the ImageNet dataset and variants of ImageNet. This demonstrates that naive fine-tuning of the entire CLIP model may cause overfitting on the training set, leading to performance degradation. MAP achieves remarkable performance on unseen data compared to zero-shot CLIP [1], linear probe CLIP, CoOp [16] and CoCoOp [18]. Compared to MaPLe, MAP shows slightly lower performance on ImageNet-Sketch but outperforms MaPLe [23] on other target datasets (ImageNetV2, ImageNet-A, and ImageNet-R). This underscores the robustness of MAP to domain shifts.
IV-D Cross-Dataset Evaluation
To demonstrate the model’s capacity for generalization beyond a single dataset, we conduct training on ImageNet [66] and subsequently evaluate its performance on the other 10 datasets. When transferring to other datasets, textual attribute prompts are constructed using class attribute descriptions of the target dataset classes, which are also collected from the LLM. The learned parameters can be directly transferred, allowing effective inference despite category differences between the source and target datasets. Table V presents a comprehensive overview of the performance comparison between MAP and previous methodologies on the cross-dataset evaluation benchmark. On the source dataset, MAP achieves the highest score, underscoring its effectiveness in the source domain. When compared with CoOp [16], CoCoOp [18], and MaPLe [23], MAP demonstrates a superior capacity for generalization across diverse datasets. Specifically, it outperforms these methodologies in 7 out of 10, 6 out of 10, and 6 out of 10 datasets, respectively. This suggests that MAP exhibits robustness to varied data distributions.
@@ IV-D Cross-Dataset Evaluation
TABLE IV: Domain generalization evaluation. Methods are trained on the source dataset ImageNet and evaluated on datasets with domain shifts, including ImageNetV2, ImageNet-S, ImageNet-A, and ImageNet-R.
@@
ImageNet ImageNetV2 ImageNet-S ImageNet-A ImageNet-R Avg.
CLIP [1] 66.73 60.83 46.15 47.77 73.96 57.18
Fully Fine-Tuned CLIP 61.65 52.70 26.10 17.55 50.15 36.63
Linear probe CLIP [1] 67.42 57.19 35.97 36.19 60.10 47.36
CoOp [16] 71.51 64.20 47.99 49.71 75.21 59.28
CoCoOp [18] 71.02 64.07 48.75 50.63 76.18 59.91
MaPLe [23] 70.72 64.07 49.15 50.90 76.98 60.27
@@
TABLE V: Cross-dataset evaluation. Models are trained on ImageNet and evaluated on target datasets. MAP achieves overall favorable performance.
@@
CoOp [16] 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55
CoCoOp [18] 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21
MaPLe [23] 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69
@@
IV-E Ablation Study
@@
Figure 6: The absolute accuracy improvements provided by using AVAE compared to scenarios without AVAE.
Figure 7: The visualization of visual attribute prompts. Guided by textual attribute semantics, visual attribute prompts focus on distinctive visual details, such as the different leaf shapes of the Moon Orchid and Japanese Anemone, the spotted coat of the Egyptian Mau, and the large ears of the Abyssinian.
Figure 8: The impact of inserting AVAE into different layers of ViT with 1 shot per class.
Effectiveness of Attribute Prompts. We denote Textual Attribute Prompts as TAP and Visual Attribute Prompts as VAP. We remove TAP and VAP from MAP as our baseline. The results in Table VI are analyzed as follows: (1) Compared to the baseline, utilizing TAP powered by the LLM effectively improves the novel accuracy, achieving an accuracy gain of 1.43%, which demonstrates textual attributes enrich the semantics for novel classes. (2) The incorporation of VAP shows a distinct performance boost on both base (+1.6%) and novel classes (+2.11%). This proves that VAP contributes to enhancing fine-grained visual perception ability by capturing visual attributes.
Figure 10: The impact of the number of textual attribute prompts per class in the base-to-novel generalization setting.
TABLE VI: Ablation results.
@@
TABLE VII: Complexity results.
CoCoOp MaPLe MAP
       parameters 0.04M 3.56M 0.74M
       GFLOPs 83.83 55.23 84.80
       test time 56.70s 9.58s 9.79s
TABLE VIII: The impact of using different LLMs.
       Method Base Novel HM
       Qwen-1.8B-Chat 97.47 73.23 83.63
       GPT-3.5 97.57 75.23 84.95
       Qwen1.5-72B-Chat 97.77 75.30 85.08
Effectiveness of Adaptive Visual Attribute Enhancement. To verify the accuracy improvement when using AVAE, we conduct few-shot image classification experiments on 6 datasets (Flowers102, DTD, UCF101, OxfordPets, Caltech101, Food101). As shown in Figure 6, the employment of AVAE brings remarkable performance gains. Furthermore, we investigate the impact of placing AVAE into different ViT layers. As observed from Figure 8, placing AVAE in the middle layers (Layer 6-8) attains superior performance. When applying AVAE in the shallow or deep layers, the performance deteriorates obviously compared to the middle layers. Therefore, the AVAE module should be placed in the middle layers. Initial visual attribute prompts can aggregate visual regional features in shallow layers and continue to capture visual attributes in the remaining layers after enhancement by AVAE.
@@
Analysis of Number of Textual Attribute Prompts. Figure 10 illustrates the averaged harmonic accuracy of using different numbers of textual attribute prompts. According to the experimental results, the introduction of textual attribute prompts indeed improves the performance, demonstrating the effectiveness of textual attribute prompts. The accuracy improves with the incorporation of more textual attribute prompts, as this introduces more descriptive information. However, when the number of textual attribute prompts exceeds four, the performance decreases. This may be attributed to the fact that additional prompts introduce more redundancy. The initial prompts are usually the most relevant and effective, while later ones may include less useful or intuitive descriptions. Increased complexity and less discriminative attributes like size or height can also burden the model, resulting in reduced performance. Overall, the accuracy changes relatively smoothly with different prompt numbers.
Impact of Different LLMs. We conduct experiments using other large language models (LLMs), specifically Qwen-1.8B-Chat and Qwen-1.5-72B-Chat [79], and examine performance variations on the Flowers102 dataset. The results in Table VIII show that Qwen-1.5-72B-Chat achieves performance comparable to GPT-3.5. However, when using Qwen-1.8B-Chat, there is a significant performance drop compared to using GPT-3.5 and Qwen-1.5-72B-Chat. This decline may be attributed to the fact that the outputs from Qwen-1.8B-Chat are sometimes inconsistent, noisy, and occasionally lack meaningful information. These findings suggest that selecting a large language model capable of generating consistent and clear outputs is crucial for maintaining performance.
Analysis of Complexity. We compare different prompting methods about the number of parameters, the GFLOPs, and the test time in Table VII. MaPLe [23] and MAP enjoy faster inference speeds than CoCoOp [18]. Compared with MaPLe, MAP is more parameter-efficient (0.74M vs 3.56M). The computation cost (GFLOPs) of MAP is higher, but considering the performance improvement, it is acceptable.
@@
V Limitation and Future Work
We use text attributes directly from GPT without manual filtering. Text attributes may contain noise that may hinder accurate classification, such as attributes with high uncertainty, like colors of toad lilies (white, purple, pink, or yellow). On Flowers102 [71], we manually filter improper attributes, resulting in an improvement of 0.37% in HM. Filtering improper ones has the potential to improve results. We’ll design an automatic filter plan in the future.
VI Conclusion
@@
Acknowledgments
This work was supported by National Defense Basic Scientific Research Program of China (JCKY2020903B002), National Natural Science Foundation of China (62306294), Anhui Provincial Natural Science Foundation (2308085QF222), China Postdoctoral Science Foundation (2023M743385) and Youth Innovation Promotion Association CAS.